OpenStudy (tiffany_rhodes):

A ball with mass m kg is thrown upward with initial velocity 19 m/s from the roof of a building 12m high. Neglect air resistance. Use g =9.8 m/s^2. Find the maximum height above the ground that the ball reaches.

hartnn (hartnn):

the max. height formula is \(\large h_{max} = \dfrac{u^2 \sin^2 \theta }{2g}\) here, since the ball is thrown vertically upwards, \(\theta = 90, \quad \sin 90 =1\) and since you're already on a 12 m height building, total height from ground H will be \(\large H= 12 +\dfrac{u^2}{2g}\) u =19 g=9.8 just plug in!

OpenStudy (tiffany_rhodes):

Thank you @hartnn!!

hartnn (hartnn):

o hey, you're welcome ^_^